OpenStudy (anonymous):

Simplify (3x + 5) + (2x - 9) - (4x + 3)

OpenStudy (pfenn1):

\[(3x + 5) + (2x - 9) - (4x + 3)\]Can you combine like terms?

OpenStudy (anonymous):

yeah, but I don't know how

OpenStudy (pfenn1):

Okay. Do you know how to distribute the negative sign outside the last term?

OpenStudy (anonymous):

no

OpenStudy (pfenn1):

Okay. By the distribution property\[-(x+y)=-x-y\]and\[-(x-y)=-x+y\]Have you seen that before?

OpenStudy (pfenn1):

So, if we distribute all the terms and get rid of the parentheses, we get\[(3x + 5) + (2x - 9) - (4x + 3)=3x+5+2x-9-4x-3\]Does that make sense?

OpenStudy (pfenn1):

Now, collect the like terms.\[(3x+2x-4x)+(5-9-3)=\]